A production filling operation has a population standard deviation of 6 ounces. When in proper adjustment, the mean filling weight for the production process is 50 ounces. A quality control inspector selects a random sample of 36 containers and finds that the mean weight is 48.6 ounces. Does the process need adjustment? Why? Find the p-value. (Hint: the process will need adjustment if the mean weight is either significantly above or below 50 ounces)
Answer by stanbon(75887)

A production filling operation has a population standard deviation of 6 ounces.
-----------
When in proper adjustment, the mean filling weight for the production process is 50 ounces.
-----------------
A quality control inspector selects a random sample of 36 containers and finds that the mean weight is 48.6 ounces.
-----------------
Does the process need adjustment? Why? Find the p-value. (Hint: the process will need adjustment if the mean weight is either significantly above or below 50 ounces)
----
Ho: u = 50 ounces
Ha: u is not equal to 50 ounces
------------------
x-bar = 48.6 ounces
---
t(48.6) = (48.6-50)/[6/sqrt(36)] = -1.4
----
p-value = 2*P(t< -1.4 when df = 35) = 0.1703
-----
Since the p-value is greater than 5%, fail to reject Ho.
The mean of the production process is statistically equal to 50;
it does not need adjustment.
==========================================================
Cheers,
Stan H.